OpenStudy (anonymous):

Find the difference quotient and simplify. f(x)=2x^2+3x-1, f(x+h)-f(x)/h, h does not equal 0

OpenStudy (anonymous):

one step at a time

OpenStudy (anonymous):

\[ f(x)=2x^2+3x-1\] \[ f(x+h)=2(x+h)^2+3(x+h)-1\] is the first step then multiply out and combine like terms

OpenStudy (anonymous):

multiply carefully \[2(x+h)^2+3(x+h)-1=2x^2+4xh+2h^2+3x+3h-1\]

OpenStudy (anonymous):

then subtract \[2x^2+4xh+2h^2+3x+3j-1-[2x^2+3x-1]\]all the terms without the \(h\) will go